# Universal mechanistic rules for de novo design of enzymes

### v1

@@ III.3.2 First passage time: speeding up a slow reaction
We now focus on the early-time dynamics of the process and quantify the characteristic time that is needed for the substrate-to-product conversion to take place, i.e. the mean first passage time. We focus on the switch reaction, although analogous results are obtained for the dissociation reaction. We initialize the system in the short state ($\phi_{\textrm{m}}=-1$) at $t=0$, and numerically solve the Langevin dynamics (Eq. (7), see Appendix C for details of the numerical solution) until the long state ($\phi_{\textrm{m}}=1$) is reached at a time $t=\tau$ (which we then record). The ensemble average of $\tau$ (i.e. average over many simulations), which we denote as $\langle\tau\rangle$, corresponds to the mean first passage time.
@@
In Fig. 8(a), we show the distribution of first passage times $P(\tau)$ for zero and large coupling strength $h$. Clearly, in the presence of dissipative coupling, the first passage times are much smaller in relative terms, while the distribution appears to be markedly narrower. Figures 8(b,c) show how the mean first passage time $\langle\tau\rangle$, relative to the value $\langle\tau_{0}\rangle$ corresponding to the absence of enzyme conformational changes ($\ell_{\textrm{e}}/\ell_{\textrm{m}}=0$), depends on the tuning parameters, namely the fuelled reaction drive $E_{*}/E_{\mathrm{ba}}$, the (geometric) deformation ratio $\ell_{\textrm{e}}/\ell_{\textrm{m}}$, and the coupling strength $h$. By comparing Fig. 8(b) with Fig. 4(a), it becomes clear that the speed-up in the reaction is directly related to the global bifurcation in the deterministic dynamics.
@@
Some analytical progress towards a calculation of the mean first passage time is possible in the limit of small noise ($k_{\rm B}T$ much smaller than the energy barriers $E_{\mathrm{ba}}$ and $\Delta$). In this limit, a generalization of the Kramers escape rate to higher dimensions due to Langer [9, 56, 57] gives the rate at which the probability current crosses through a saddle point, which represents the transition state. To deal with the multiplicative nature of the noise, we evaluate the mobility tensor $M_{\alpha\beta}$ around the saddle point of interest. Moreover, we are interested in the catalyzed reaction, and thus focus on the transition state shown as the red diamond in Fig. 3. The rate then takes a similar form to the one-dimensional case (see Eq. (11)), given by
$k_{\rm Langer}=\frac{|\Lambda_{-}|}{2\pi}\sqrt{\frac{\lambda_{\mathrm{e}}^{(0)}}{|\lambda_{\mathrm{e}}|}}\exp\bigg{(}{-\frac{E_{\mathrm{ba}}}{k_{\rm B}T}}\bigg{)}.$ (16)
@@
For more details see Appendix E. As shown in Fig. 8(d), Eq. (16) agrees well with numerical solution of Eq. (5) for the crossing rate over the saddle point marked as the red diamond in Fig. 3, after a brief initial transient. In the numerics, the initial probability distribution for both enzyme and molecule was chosen as a narrow Gaussian centered at the potential minimum.
@@ IV Discussion
We have developed a minimal model for a fuelled enzyme that is able to extract energy from a thermodynamically favourable reaction to drive and speed-up a thermodynamically unfavourable reaction. The transduction is mechanical, arising from the conformational changes of the enzyme that couple to the conformational changes of the attached substrate molecule. This goes beyond previously developed minimal models for enzymes that only described passive enzymes [36, 37, 38, 39, 40, 41], which can speed up energetically favourable reactions but are unable to alter the reaction equilibrium as determined by thermodynamics.
Figure 7: (a) Probability of finding the shape-switching molecule in the long state and (b) energy dissipation rate, versus the length ratio $\ell_{\textrm{e}}/\ell_{\textrm{m}}$ and the coupling strength $h$. The black line corresponds to $P_{\mathrm{long}}=0.5$. The noise strength is $k_{\rm B}T/\Delta=0.2$.
@@
Figure 8: (a) Distribution of first passage times $\tau$ for $\ell_{\textrm{e}}/\ell_{\textrm{m}}=1.6$ in the absence (red) and presence (green) of dissipative coupling. (b,c) Mean first passage time $\langle\tau\rangle$, normalized by the mean first passage time in the uncoupled case where $\ell_{\textrm{e}}/\ell_{\textrm{m}}=0$, as a function of the length ratio $\ell_{\textrm{e}}/\ell_{\textrm{m}}$ and the energy ratio $E_{*}/E_{\mathrm{ba}}$ for fixed value of coupling strength $h=0.48$ (a), and as a function of the coupling strength $h$ and $E_{*}/E_{\mathrm{ba}}$ for fixed $\ell_{\textrm{e}}/\ell_{\textrm{m}}=1$ in (b). In (a–c), the noise strength is $k_{\rm B}T/\Delta=0.15$. (d) The rate at which the saddle point marked as the red diamond in Fig. 3 is crossed extracted numerically (coloured lines) compared to the Langer rate given by Eq. (16) (black horizontal lines), for $E_{\mathrm{ba}}/\Delta=0.8$ and $k_{\rm B}T/\Delta=0.08$.
The function of our model enzyme is most prominently dictated by the geometry of the enzyme-substrate complex, and can be optimized by following three simple golden rules: (i) the enzyme and the molecule should be attached at the smaller end of each; (ii) the conformational change of the enzyme must be comparable to or larger than the conformational change required of the molecule; (iii) the conformational change of the enzyme must be fast enough. These parameters can be experimentally tuned when designing an artificial enzyme, and should be experimentally accessible when investigating the function of biological enzymes. Interestingly, the effectiveness of the enzyme (as measured by successful catalytic action) and its energy dissipation do not follow a one-to-one relation (Fig. 7), leaving additional room for the optimization of not just the enzyme effectiveness but its energetic efficiency.
@@
We were able to understand the emergence of a strong coupling between the thermodynamically favourable and unfavourable reactions to be a result of a global bifurcation in the deterministic dynamics of the system, of a similar nature to that causing synchronization and phase-locking among mechanically-coupled enzymes [35, 47]. Importantly, however, by adding thermodynamically-consistent fluctuations to the theory, we showed that the effect survives in the presence of noise, and we could calculate important quantities such as non-equilibrium steady states reflecting the departure from equilibrium thermodynamics in the reaction equilibrium, and first passage time distributions for the catalyzed reaction.
Our approach readily allows us to probe the energetic aspects of the non-equilibrium reaction [58], and in particular, the entropy production due to the coupling between the mechanical and the chemical degrees of freedom [59]. It can also be used to study evolutionary advantages of having mechanical control during the catalytic process via conformational changes, as opposed to simple utilization of the uncontrolled metal-based catalysis.
@@ Acknowledgements.
We acknowledge discussions with Navdeep Rana regarding the numerical integrators. We acknowledge support from the Max Planck School Matter to Life and the MaxSynBio Consortium which are jointly funded by the Federal Ministry of Education and Research (BMBF) of Germany and the Max Planck Society.
Appendix A Derivation of the phase-equations
@@
$\dot{L}_{\textrm{e}}=-\mu_{1}k[L_{\textrm{e}}-L(\phi_{\textrm{e}})]+\mu_{\textrm{b}}\partial_{L_{\textrm{m}}}V_{\textrm{m}}(L_{\textrm{m}}),$ (17)
$\dot{L}_{\textrm{m}}=\mu_{\textrm{b}}k[L_{\textrm{e}}-L(\phi_{\textrm{e}})]-\mu_{2}\partial_{L_{\textrm{m}}}V_{\textrm{m}}(L_{\textrm{m}}).$ (18)
@@
$\dot{\phi}_{\textrm{e}}$ $=\mu_{\phi}[-\partial_{\phi_{\textrm{e}}}U(L_{\textrm{e}},\phi_{\textrm{e}})],$ (19)
@@
with the prime representing a derivative with respect to $\phi_{\textrm{e}}$. We define $\delta L_{\textrm{e}}\equiv L_{\textrm{e}}-L(\phi_{\textrm{e}})$ such that $\dot{L}_{\textrm{e}}=\dot{\delta L_{\textrm{e}}}+L^{\prime}(\phi_{\textrm{e}})\dot{\phi}_{\textrm{e}}$. If the enzyme is relatively stiff with the timescale $(\mu_{1}k)^{-1}$ being much shorter than the timescale of changes in the internal phase, the enzyme length will quickly adapt to changes of the preferred length, and thus $\dot{\delta L_{\textrm{e}}}\approx 0$. From Eq. (17), this implies
$L^{\prime}(\phi_{\textrm{e}})\dot{\phi}_{\textrm{e}}\approx-\mu_{1}k[L_{\textrm{e}}-L(\phi_{\textrm{e}})]+\mu_{\textrm{b}}\partial_{L_{\textrm{m}}}V(L_{\textrm{m}}),$ (20)
@@
which allows us to eliminate $k[L_{\textrm{e}}-L(\phi_{\textrm{e}})]$ between Eqs. (18) and (A), hence projecting the dynamics onto the slow manifold $(\phi_{\textrm{e}},L_{\textrm{m}})$. Finally, we reexpress $L_{\textrm{m}}$ in terms of the dimensionless reaction coordinate $\phi_{\textrm{m}}$ through $L_{\textrm{m}}=L_{\textrm{m}}^{(0)}+\ell_{\textrm{m}}\phi_{\textrm{m}}$. The resulting governing equations for $(\phi_{\textrm{e}},\phi_{\textrm{m}})$ then reduce to Eq. (1).
Appendix B Reverse switch reaction
@@
One can observe that for the switch reaction our enzyme is equally capable of catalyzing a reverse reaction, i.e. from long to short conformational state, by choosing $\varepsilon<0$. Exploiting the symmetries of our dynamical equations, we find that a $\pi$ phase shift will be needed for $\phi_{\textrm{e}}$ in the conformation cycle (Eq. (8)), as in this case the enzyme will first contract and then expand during its conformational change. A phase portrait of the deterministic dynamics at strong coupling (beyond the global bifurcation) is shown in Fig. 9(a). This can be compared to the dynamics of the forward switch reaction in Fig. 3(d). Note that the bifurcation now happens near the upper saddle point, where the blue solid line represents the flow from the long state to the short state. An example of the NESS distributions for this reverse case is shown in Fig. 9(b), which may be compared to the case of the forward reaction in Fig. 5(b). It highlights how the action of the enzyme can favour the short state even if the long state is thermodynamically preferred.
Appendix C Simulations
To generate the phase portraits in Fig. 3 and the parameter scan in Fig. 4 from the deterministic equations (Eq. (1)), we used the built-in ode45 integrator of MATLAB, which uses a fourth order Runge-Kutta method [60]. A grid of $201\times 201$ for $-\pi<\phi_{\textrm{e}}<\pi$ and $-2<\phi_{\textrm{m}}<2$ was used to generate the initial conditions, and the equations were integrated for a total time of $\tilde{t}_{\rm tot}=120$.
For the numerical integration of the Fokker-Planck equation given by Eq. (5) we used a custom code written in Python [61]. The code uses a finite difference method with fourth-order accuracy for the spatial derivatives and fourth order Runge-Kutta method for the time integration. For the enzyme coordinate $\phi_{\textrm{e}}$ we considered the domain $0\leq\phi_{\textrm{e}}<2\pi$ with periodic boundary conditions. For the molecule coordinate $\phi_{\textrm{m}}$, we used Dirichlet boundary conditions in the domains $[-1.8,1.8],[-2.0,2.0],[-2.2,2.2]$ and noise strengths $k_{\rm B}T/\Delta=0.15,0.2,0.4$. To discretize the domains we used a rectangular shape domain such that the finite-difference increments satisfied $\Delta\phi_{\textrm{e}}\simeq\Delta\phi_{\textrm{m}}$ and thus the integrator would be more stable. For $k_{\rm B}T/\Delta=0.15,0.2,0.4$ the corresponding number of points $(N_{1},N_{2})$ per side were $(512,128),(400,128),(300,100)$, and the time step was $\Delta\tilde{t}=10^{-3},10^{-3},5\cdot 10^{-4}$.
For the numerical integration of the Langevin equations in Eq. (7), we used a custom code written in Julia [62]. The Euler-Maruyama method was used for the numerical integration. The time step used was $\Delta\tilde{t}=10^{-3}$.
Figure 9: Reverse switch reaction. (a) Phase-portrait for $h=0.48$, $\ell_{\textrm{e}}/\ell_{\textrm{m}}=2.5$, and $\varepsilon/\Delta=-0.1$. The black solid line represents a deterministic trajectory. (b) Marginal distributions of the molecule reaction coordinate in steady-state, for various values of the coupling strength. (a) and (b) can be respectively compared to Fig. 3(d) and Fig. 5(b) for the forward reaction.
To check for consistency between the Langevin and Fokker-Planck approaches, long trajectories from the solution of the Langevin equations were binned and compared with the steady state distribution predicted by the Fokker-Planck equation. The two approaches exhibit good agreement, as depicted in Fig. 10 for the case of $h=0.67$ and $k_{\rm B}T/\Delta=0.2$.
Appendix D Steady state distribution for an uncoupled enzyme
Figure 10: Comparison of the marginal distributions obtained from the Fokker-Planck equation (red solid line) at steady state and long simulations of the Langevin equations (histogram). (a) and (b) show the distribution of the enzyme and molecule reaction coordinates, respectively. The parameters for this example are $h=0.67$ and $k_{\rm B}T/\Delta=0.2$.
@@ Appendix D Steady state distribution for an uncoupled enzyme
$J_{\mathrm{e}}=-M_{\mathrm{ee}}(\phi_{\textrm{e}})[V^{\prime}_{\textrm{e}}(\phi_{\textrm{e}})P_{\mathrm{e}}+k_{\rm B}T\partial_{\phi_{\textrm{e}}}P_{\mathrm{e}}].$ (21)
@@
$+\int^{\phi_{\textrm{e}}}_{0}dx\,M_{\textrm{e}\textrm{e}}^{-1}(x)\exp\bigg{(}\frac{V_{\textrm{e}}(x)-V_{\textrm{e}}(\phi_{\textrm{e}})-2\pi F}{k_{\rm B}T}\bigg{)}\bigg{]}.$ (24)
The form of $V_{\textrm{e}}$ does not readily lend itself to further analytical progress, and we integrate Eq. (D) numerically to obtain the black lines in Fig. 5(a,c). In the particular case $E_{*}=0$, we find that $J_{\textrm{e}}=0$ and thus the enzyme coordinate is in thermodynamic equilibrium described by the Boltzmann weight $P_{\textrm{e}}(\phi_{\textrm{e}})\propto e^{-V_{\textrm{e}}(\phi_{\textrm{e}})/k_{\rm B}T}$.
Appendix E Langer rate
@@
$k_{\rm Langer}=\frac{|\Lambda_{-}|}{2\pi}\frac{\sqrt{\lambda_{\mathrm{e}}^{(0)}\lambda_{\mathrm{m}}^{(0)}}}{\sqrt{|\lambda_{\mathrm{e}}|\lambda_{\mathrm{m}}}}\exp\bigg{(}{-\frac{E_{\mathrm{ba}}}{k_{\rm B}T}}\bigg{)},$ (25)
where $\lambda_{\alpha}^{(0)}=V^{\prime\prime}_{\alpha}(\phi_{\alpha})|_{\rm min}$, and $\lambda_{\alpha}=V^{\prime\prime}_{\alpha}(\phi_{\alpha})|_{\rm saddle}$, namely, the corresponding values evaluated around the saddle point. The definitions imply that $\lambda_{\mathrm{m}}=\lambda_{\mathrm{m}}^{(0)}$, and thus, they cancel each other in Eq. (25). Finally, $\Lambda_{-}$ is the negative eigenvalue of the matrix $\lambda_{\alpha}M_{\alpha\beta}$ evaluated around the saddle point which is given by
$\Lambda_{-}=\frac{1}{2}\bigg{(}M_{\textrm{e}\textrm{e}}\lambda_{\textrm{e}}+M_{\textrm{m}\textrm{m}}\lambda_{\textrm{m}}$
$\hskip 14.22636pt-\sqrt{(M_{\textrm{e}\textrm{e}}\lambda_{e}-M_{\textrm{m}\textrm{m}}\lambda_{\textrm{m}})^{2}+4M_{\textrm{e}\textrm{m}}^{2}\lambda_{\textrm{e}}\lambda_{\textrm{m}}}\bigg{)}.$ (26)

### v2

@@ III.3.2 First passage time: speeding up a slow reaction
We now focus on the early-time dynamics of the process and quantify the characteristic time that is needed for the substrate-to-product conversion to take place, i.e. the mean first passage time. We focus on the switch reaction, although analogous results are obtained for the dissociation reaction. We initialize the system in the short state ($\phi_{\textrm{m}}=-1$) at $t=0$, and numerically solve the Langevin dynamics (Equation (7), see Methods for details of the numerical solution) until the long state ($\phi_{\textrm{m}}=1$) is reached at a time $t=\tau_{\mathrm{fpt}}$ (which we then record). The ensemble average of $\tau_{\mathrm{fpt}}$ (i.e. average over many simulations), which we denote as $\langle\tau_{\mathrm{fpt}}\rangle$, corresponds to the mean first passage time.
@@
In Figure 8(A), we show the distribution of first passage times $P(\tau_{\mathrm{fpt}})$ for zero and large coupling strength $h$. Clearly, in the presence of dissipative coupling, the first passage times are much smaller in relative terms, while the distribution appears to be markedly narrower. Figures 8(B,C) show how the mean first passage time $\langle\tau_{\mathrm{fpt}}\rangle$, relative to the value $\langle\tau_{\mathrm{fpt,0}}\rangle$ corresponding to the absence of enzyme conformational changes ($\ell_{\textrm{e}}/\ell_{\textrm{m}}=0$), depends on the tuning parameters, namely the fuelled reaction drive $E_{*}/E_{\mathrm{ba}}$, the (geometric) deformation ratio $\ell_{\textrm{e}}/\ell_{\textrm{m}}$, and the coupling strength $h$. By comparing Figure 8(B) with Figure 4(A), it becomes clear that the speed-up in the reaction is directly related to the global bifurcation in the deterministic dynamics.
@@
Some analytical progress towards a calculation of the mean first passage time is possible in the limit of small noise ($k_{\rm B}T$ much smaller than the energy barriers $E_{\mathrm{ba}}$ and $\Delta$). In this limit, a generalization of the Kramers escape rate to higher dimensions due to Langer [9, 59, 60] gives the rate at which the probability current crosses through a saddle point, which represents the transition state. To deal with the multiplicative nature of the noise, we evaluate the mobility tensor $M_{\alpha\beta}$ around the saddle point of interest. Moreover, we are interested in the catalyzed reaction, and thus focus on the transition state shown as the red diamond in Figure 3. The rate then takes a similar form to the one-dimensional case (see Equation (11)), given by
$k(\mathrm{fs}\to\mathrm{wp})=\frac{|\Lambda_{-}|}{2\pi}\sqrt{\frac{\lambda_{\mathrm{e}}^{(0)}}{|\lambda_{\mathrm{e}}^{(1)}|}}\exp\bigg{(}{-\frac{E_{\mathrm{ba}}}{k_{\rm B}T}}\bigg{)}$ (16)
@@
where $\lambda_{\mathrm{e}}^{(0)}=V^{\prime\prime}_{\mathrm{e}}(\phi_{\mathrm{e}})|_{\rm min}$, $\lambda_{\mathrm{e}}^{(1)}=V^{\prime\prime}_{\mathrm{e}}(\phi_{\mathrm{e}})|_{\rm saddle}$, and $\Lambda_{-}$ along with additional details is given in the Supplemental Information. As shown in Figure 8(D), Equation (16) agrees well with numerical solution of Equation (5) for the crossing rate over the saddle point marked as the red diamond in Figure 3, after a brief initial transient. In the numerics, the initial probability distribution for both enzyme and molecule was chosen as a narrow Gaussian centered at the potential minimum.
@@ IV Discussion
We have developed a minimal model for a fuelled enzyme that is able to extract energy from a thermodynamically favourable reaction to drive and speed-up a thermodynamically unfavourable reaction. Hence, the enzyme acts as a “Maxwell’s demon”. The transduction is mechanical, arising from the conformational changes of the enzyme that couple to the conformational changes of the attached substrate molecule. This goes beyond previously developed minimal models for enzymes that only described passive enzymes [37, 38, 39, 40, 42, 43], which can speed up energetically favourable reactions but are unable to alter the reaction equilibrium as determined by thermodynamics.
@@
Figure 8: Results on first passage times (A) Distribution of first passage times $\tau_{\mathrm{fpt}}$ for $\ell_{\textrm{e}}/\ell_{\textrm{m}}=1.6$ in the absence (red) and presence (green) of dissipative coupling. (B,C) Mean first passage time $\langle\tau_{\mathrm{fpt}}\rangle$, normalized by the mean first passage time in the uncoupled case where $\ell_{\textrm{e}}/\ell_{\textrm{m}}=0$, as a function of the length ratio $\ell_{\textrm{e}}/\ell_{\textrm{m}}$ and the energy ratio $E_{*}/E_{\mathrm{ba}}$ for fixed value of coupling strength $h=0.48$ (A), and as a function of the coupling strength $h$ and $E_{*}/E_{\mathrm{ba}}$ for fixed $\ell_{\textrm{e}}/\ell_{\textrm{m}}=1$ in (B). In (A–C), the noise strength is $k_{\rm B}T/\Delta=0.15$. (D) The rate at which the saddle point marked as the red diamond in Figure 3 is crossed extracted numerically (coloured lines) compared to the Langer rate given by Equation (16) (black horizontal lines), for $E_{\mathrm{ba}}/\Delta=0.8$ and $k_{\rm B}T/\Delta=0.08$.
[…]
@@
In general, we were able to understand the emergence of a strong coupling between the thermodynamically favourable and unfavourable reactions to be a result of a global bifurcation in the deterministic dynamics of the system, of a similar nature to that causing synchronization and phase-locking among mechanically-coupled enzymes [36, 50], which can explain the enhanced activity of enzyme oligomeric complexes as observed in some experiments [51, 52]. Importantly, by adding thermodynamically-consistent fluctuations to the theory, we showed that the effect survives in the presence of noise, and we could calculate important quantities such as non-equilibrium steady states reflecting the departure from equilibrium thermodynamics in the reaction equilibrium, and first passage time distributions for the catalyzed reaction. Interestingly, the effectiveness of the enzyme (as measured by successful catalytic action) and its energy dissipation do not follow a one-to-one relation (Figure 7), leaving additional room for the optimization of not just the enzyme effectiveness but its energetic efficiency.
Our approach readily allows us to probe the energetic aspects of the non-equilibrium reaction [68], and in particular, the entropy production due to the coupling between the mechanical and the chemical degrees of freedom [69]. The proposed design rules in this work focused on optimizing the transition step of the reaction, assuming that the binding of the substrate and unbinding of the product from the enzyme is essentially instantaneous and perfectly specific, and that the enzyme-substrate bond is strong enough to avoid unbinding of the substrate during the (incomplete) reaction. A future and more detailed description of these binding and unbinding processes is an obvious next step for the further optimization of the enzymatic activity, now including selectivity. In any case, the present mechanistic framework (and the resulting rules) can provide physical intuition to be used as input to complement atomistic simulation and machine learning approaches for de novo design of proteins with dynamic function such as enzymes [70]. It can also be used to study the evolutionary advantage of having mechanical control during the catalytic process via conformational changes, as opposed to simple utilization of uncontrolled metal-based catalysis.
@@ Acknowledgements.
We acknowledge discussions with Navdeep Rana regarding the numerical integrators, and support from the Max Planck School Matter to Life and the MaxSynBio Consortium which are jointly funded by the Federal Ministry of Education and Research (BMBF) of Germany and the Max Planck Society.
Methods
Simulations
To generate the phase portraits in Figure 3 and the parameter scan in Figure 4 from the deterministic equations (Equation (1)), we used the built-in ode45 integrator of MATLAB, which uses a fourth order Runge-Kutta method [71]. A grid of $201\times 201$ for $-\pi<\phi_{\textrm{e}}<\pi$ and $-2<\phi_{\textrm{m}}<2$ was used to generate the initial conditions, and the equations were integrated for a total time of $\tilde{t}_{\rm tot}=120$.
For the numerical integration of the Fokker-Planck equation given by Equation (5) we used a custom code written in Python [72]. The code uses a finite difference method with fourth-order accuracy for the spatial derivatives and fourth order Runge-Kutta method for the time integration. For the enzyme coordinate $\phi_{\textrm{e}}$ we considered the domain $0\leq\phi_{\textrm{e}}<2\pi$ with periodic boundary conditions. For the molecule coordinate $\phi_{\textrm{m}}$, we used Dirichlet boundary conditions in the domains $[-1.8,1.8],[-2.0,2.0],[-2.2,2.2]$ and noise strengths $k_{\rm B}T/\Delta=0.15,0.2,0.4$. To discretize the domains we used a rectangular shape domain such that the finite-difference increments satisfied $\Delta\phi_{\textrm{e}}\simeq\Delta\phi_{\textrm{m}}$ and thus the integrator would be more stable. For $k_{\rm B}T/\Delta=0.15,0.2,0.4$ the corresponding number of points $(N_{1},N_{2})$ per side were $(512,128),(400,128),(300,100)$, and the time step was $\Delta\tilde{t}=10^{-3},10^{-3},5\cdot 10^{-4}$.
For the numerical integration of the Langevin equations in Equation (7), we used a custom code written in Julia [73]. The Euler-Maruyama method was used for the numerical integration. The time step used was $\Delta\tilde{t}=10^{-3}$.
To check for consistency between the Langevin and Fokker-Planck approaches, long trajectories from the solution of the Langevin equations were binned and compared with the steady state distribution predicted by the Fokker-Planck equation. The two approaches exhibit good agreement, as depicted in Figure S2 for the case of $h=0.67$ and $k_{\rm B}T/\Delta=0.2$.
Supplemental Information
Michalis Chatzittofi Ramin Golestanian Jaime Agudo-Canalejo
May 14, 2025
Supplemental Methods
Derivation of the phase-equations
@@ Methods
$\dot{L}_{\textrm{e}}=-\mu_{1}k[L_{\textrm{e}}-L(\phi_{\textrm{e}})]+\mu_{\textrm{b}}\partial_{L_{\textrm{m}}}V_{\textrm{m}}(L_{\textrm{m}}),$ (S1)
$\dot{L}_{\textrm{m}}=\mu_{\textrm{b}}k[L_{\textrm{e}}-L(\phi_{\textrm{e}})]-\mu_{2}\partial_{L_{\textrm{m}}}V_{\textrm{m}}(L_{\textrm{m}}).$ (S2)
@@
$\dot{\phi}_{\textrm{e}}$ $=\mu_{\phi}[-\partial_{\phi_{\textrm{e}}}U(L_{\textrm{e}},\phi_{\textrm{e}})],$ (S3)
@@
with the prime representing a derivative with respect to $\phi_{\textrm{e}}$. We define $\delta L_{\textrm{e}}\equiv L_{\textrm{e}}-L(\phi_{\textrm{e}})$ such that $\dot{L}_{\textrm{e}}=\dot{\delta L_{\textrm{e}}}+L^{\prime}(\phi_{\textrm{e}})\dot{\phi}_{\textrm{e}}$. If the enzyme is relatively stiff with the timescale $(\mu_{1}k)^{-1}$ being much shorter than the timescale of changes in the internal phase, the enzyme length will quickly adapt to changes of the preferred length, and thus $\dot{\delta L_{\textrm{e}}}\approx 0$. From Equation (S1), this implies
$L^{\prime}(\phi_{\textrm{e}})\dot{\phi}_{\textrm{e}}\approx-\mu_{1}k[L_{\textrm{e}}-L(\phi_{\textrm{e}})]+\mu_{\textrm{b}}\partial_{L_{\textrm{m}}}V(L_{\textrm{m}}),$ (S4)
or equivalently
$k[L_{\textrm{e}}-L(\phi_{\textrm{e}})]\approx\frac{1}{\mu_{1}}\left(\mu_{\textrm{b}}\partial_{L_{\textrm{m}}}V(L_{\textrm{m}})-L^{\prime}(\phi_{\textrm{e}})\dot{\phi}_{\textrm{e}}\right).$ (S5)
@@
Substitution of Equation (S5) into Equations  (S2) and (S3) allows us to eliminate $k[L_{\textrm{e}}-L(\phi_{\textrm{e}})]$ from these equations, which we then solve for $\dot{\phi}_{\textrm{e}}$ and $\dot{L}_{\textrm{m}}$. In this way, we project the dynamics onto the slow manifold $(\phi_{\textrm{e}},L_{\textrm{m}})$. Finally, we reexpress $L_{\textrm{m}}$ in terms of the dimensionless reaction coordinate $\phi_{\textrm{m}}$ through $L_{\textrm{m}}=L_{\textrm{m}}^{(0)}+\ell_{\textrm{m}}\phi_{\textrm{m}}$. The resulting governing equations for $(\phi_{\textrm{e}},\phi_{\textrm{m}})$ then correspond to Equation (1) in the main article.
Reverse switch reaction
@@
One can observe that for the switch reaction our enzyme is equally capable of catalyzing a reverse reaction, i.e. from long to short conformational state, by choosing $\varepsilon<0$. Exploiting the symmetries of our dynamical equations, we find that a $\pi$ phase shift will be needed for $\phi_{\textrm{e}}$ in the conformation cycle (Equation 8), as in this case the enzyme will first contract and then expand during its conformational change. A phase portrait of the deterministic dynamics at strong coupling (beyond the global bifurcation) is shown in Figure S1(A). This can be compared to the dynamics of the forward switch reaction in Figure 3(D). Note that the bifurcation now happens near the upper saddle point, where the blue solid line represents the flow from the long state to the short state. An example of the NESS distributions for this reverse case is shown in Figure S1(B), which may be compared to the case of the forward reaction in Figure 5(B). It highlights how the action of the enzyme can favour the short state even if the long state is thermodynamically preferred.
Steady state distribution for an uncoupled enzyme
@@
$J_{\mathrm{e}}=-M_{\mathrm{ee}}(\phi_{\textrm{e}})[V^{\prime}_{\textrm{e}}(\phi_{\textrm{e}})P_{\mathrm{e}}+k_{\rm B}T\partial_{\phi_{\textrm{e}}}P_{\mathrm{e}}].$ (S6)
@@
$+\int^{\phi_{\textrm{e}}}_{0}dx\,M_{\textrm{e}\textrm{e}}^{-1}(x)\exp\bigg{(}\frac{V_{\textrm{e}}(x)-V_{\textrm{e}}(\phi_{\textrm{e}})-2\pi F}{k_{\rm B}T}\bigg{)}\bigg{]}.$ (S9)
The form of $V_{\textrm{e}}$ does not readily lend itself to further analytical progress, and we integrate Equation (Steady state distribution for an uncoupled enzyme) numerically to obtain the black lines in Figure 5(A,B). In the particular case $E_{*}=0$, we find that $J_{\textrm{e}}=0$ and thus the enzyme coordinate is in thermodynamic equilibrium described by the Boltzmann weight $P_{\textrm{e}}(\phi_{\textrm{e}})\propto e^{-V_{\textrm{e}}(\phi_{\textrm{e}})/k_{\rm B}T}$.
Langer rate
@@
$k(\mathrm{fs}\to\mathrm{wp})=\frac{|\Lambda_{-}|}{2\pi}\frac{\sqrt{\lambda_{\mathrm{e}}^{(0)}\lambda_{\mathrm{m}}^{(0)}}}{\sqrt{|\lambda_{\mathrm{e}}^{(1)}|\lambda_{\mathrm{m}}^{(1)}}}\exp\bigg{(}{-\frac{E_{\mathrm{ba}}}{k_{\rm B}T}}\bigg{)},$ (S10)
where $\lambda_{\alpha}^{(0)}=V^{\prime\prime}_{\alpha}(\phi_{\alpha})|_{\rm min}$, and $\lambda_{\alpha}^{(1)}=V^{\prime\prime}_{\alpha}(\phi_{\alpha})|_{\rm saddle}$, namely, the corresponding values evaluated around the saddle point. The definitions imply that $\lambda_{\mathrm{m}}^{(1)}=\lambda_{\mathrm{m}}^{(0)}$, and thus, they cancel each other in Equation (S10). Finally, $\Lambda_{-}$ is the negative eigenvalue of the matrix $\lambda_{\alpha}^{(1)}M_{\alpha\beta}$ evaluated around the saddle point which is given by
$\Lambda_{-}=\frac{1}{2}\bigg{(}M_{\textrm{e}\textrm{e}}\lambda_{\textrm{e}}^{(1)}+M_{\textrm{m}\textrm{m}}\lambda_{\textrm{m}}^{(1)}$
$\hskip 14.22636pt-\sqrt{(M_{\textrm{e}\textrm{e}}\lambda_{\mathrm{e}}^{(1)}-M_{\textrm{m}\textrm{m}}\lambda_{\textrm{m}}^{(1)})^{2}+4M_{\textrm{e}\textrm{m}}^{2}\lambda_{\textrm{e}}^{(1)}\lambda_{\textrm{m}}^{(1)}}\bigg{)}.$ (S11)
Supplemental Figures
Supplementary Fig. S1: Reverse switch reaction. (A) Phase-portrait for $h=0.48$, $\ell_{\textrm{e}}/\ell_{\textrm{m}}=2.5$, and $\varepsilon/\Delta=-0.1$. The black solid line represents a deterministic trajectory. (B) Marginal distributions of the molecule reaction coordinate in steady-state, for various values of the coupling strength. (A) and (B) can be respectively compared to Figure 3(D) and Figure 5(C) for the forward reaction.
Supplementary Fig. S2: Comparison of the marginal distributions obtained from the Fokker-Planck equation at steady state and long simulations of the Langevin equations. (A) and (B) show the distribution of the enzyme and molecule reaction coordinates, respectively. The parameters for this example are $h=0.67$ and $k_{\rm B}T/\Delta=0.2$. The red solid line is the solution of the Fokker-Planck equation and the histograms are obtained by the Langevin equations.